# Quantum Implicit Neural Representations

### v2

@@ 2.2 Data Re-uploading Quantum Circuits
Pérez-Salinas et al. (2020) first introduced the concept of data re-uploading quantum circuits and demonstrated their capability in handling certain point set classification tasks. Wach et al. (2023) further explored their classification performance and conducted experiments on image classification. Some studies concentrate on investigating the mathematical properties of data re-uploading quantum circuits. Schuld et al. (2021) revealed the fundamental nature of data re-uploading quantum circuits as Fourier series. Yu et al. (2022) further investigated the universality of single-qubit data re-uploading circuits. Shin et al. (2023) utilized a preprocessing method to enhance the ability of data re-uploading quantum circuits to fit Fourier series, but this method lacks flexibility and is difficult to apply in practice. The aforementioned studies have only experimented on simple synthetic datasets and have not found a suitable practical application scenario for data re-uploading quantum circuits.
@@ 4.3 Theoretical Analysis of QIREN
By observing Eq. (13), we can understand that applying a Linear layer before uploading the data to the quantum circuit is equivalent to adjusting the eigenvalues in the Hamiltonian, ultimately affecting the frequency spectrum. Using the Linear layer brings two advantages. First, it can make the frequency spectrum larger. From Eq. (18) and (19), it can be deduced that certain sums of eigenvalues generate the same term in the spectrum. This redundancy can be reduced by using the Linear layer, and the size of the spectrum can be extended from $(2dL+1)^{d_{h}}$ to up to $((3^{d}-1)L+1)^{d_{h}}$. Second, it enables the coverage range of the frequency spectrum to be adjustable, aiming to cover frequencies with larger coefficients, which are more important. Therefore, incorporating Linear layers can further enhance the fitting performance of QIREN. A more detailed explanation of Claim 2 can be found in Appendix B.
@@ Appendix B Additional Explanation on Claim 2
Firstly, we introduce an exponential encoding scheme (Shin et al., 2023), where the specific form of $H$ is as follows:
$H=\sum_{q=1}^{d}\beta_{q}Z^{(q)}/2.$ (A3)
In the absence of degeneracy, there can be $2^{d}$ distinct eigenvalues. Thus, the spectrum can contain at most $14^{d}-2^{d}+$ 1 distinct frequencies (including zero). But if we naively encode all $\beta_{q}=1$, then there will be only $d+1$ distinct eigenvalues and the size of the spectrum $\Omega=\left\{\lambda_{k}-\lambda_{j}\right\}$ will be limited to $12d+$ 1 . Therefore, Shin et al. (2023) derived how to choose the appropriate $\beta$ values to maximize the size of the spectrum.
@@
With $H=\beta_{1}Z/2$ and upon setting $\beta_{1}=1$, we have $\Omega^{(1)}=\{-1,0,1\}$ as the frequency spectrum. If we append one more encoding gate to $H$,we have $\Omega^{(2)}=\{-\beta_{2}-1,-\beta_{2},-\beta_{2}+1,-1,0,1,\beta_{2}-1,\beta_{2},\beta_{2}+1\}$. More generally, let $\Omega^{(k)}$ as the $k$-th frequency spectrum as a result of using $k$ encoding gates. Then $\Omega^{(k)}$ has all elements from $\Omega^{(k-1)}$, along with the new ones generated by adding ±$\pm\beta_{k}$ to all elements of $\Omega^{(k-1)}$:
$\Omega^{(k)}=\left\{\Omega^{(k-1)}-\beta_{k},\Omega^{(k-1)},\Omega^{(k-1)}+\beta_{k}\right\}.$ (A4)
@@
As $\Omega^{(k)}$ is always symmetric about zero, to generate a maximally non-degenerate integer-valued frequency spectrum, the inequality
$\max\left\{\alpha\in\Omega^{(k-1)}\right\}<\beta_{k}-\max\left\{\alpha\in\Omega^{(k-1)}\right\}$ (A5)
@@
is to be satisfied, where we can easily deduce that $\max\left\{\alpha\in\Omega^{(k-1)}\right\}=\sum_{j=1}^{k-1}\beta_{j}$. To construct a dense integer-valued spectrum starting from zero, the following recursive equation
$2\sum_{j=1}^{k-1}\beta_{j}+1=\beta_{k}$ (A6)
must be satisfied. When $\beta_{1}=1$, we have $\beta_{k}=3^{k-1}$, which is the exponential encoding scheme. We can also construct a non-dense spectrum by choosing large spacings. By applying this scheme, the size of the spectrum is extended from $12d+$ 1 to $3^{d}$.
Next, we analyze the Linear layer before data re-uploading circuit. If we ignore the bias term, the Linear layer essentially maps the input $x$ to $(w_{1}x,\ldots,w_{d}x)$. Considering the form of the encoding gate as $S(x)=\otimes_{q=1}^{d}\mathrm{e}^{-\mathrm{i}(w_{q}x)Z/2}=\mathrm{e}^{-\mathrm{i}x(w_{q}Z/2)}$, it can be concluded that the Linear layer can be equivalent to an encoding scheme. This encoding scheme is learned through training, and the upper bound is to extend the size of the spectrum to $3^{d}$ just as the exponential encoding scheme does. In addition, unlike exponential encoding scheme, the Linear layer encoding scheme are more flexible, with advantages such as not being limited to integer spectrum and the sparsity of the spectrum being learnable and adaptive. This allows us to realize an end-to-end training process without manually adjusting the hyperparameters $\beta$ for different signals.

### v3

@@ 2.2 Data Re-uploading Quantum Circuits
Pérez-Salinas et al. (2020) first introduced the concept of data re-uploading quantum circuits and demonstrated their capability in handling certain point set classification tasks. Wach et al. (2023) further explored their classification performance and conducted experiments on image classification. Some studies concentrate on investigating the mathematical properties of data re-uploading quantum circuits. Schuld et al. (2021) revealed the fundamental nature of data re-uploading quantum circuits as Fourier series. Yu et al. (2022) further investigated the universality of single-qubit data re-uploading circuits. These works only explore the multi-layer single-qubit circuits or single-layer multi-qubit circuits. We extended our research to multi-layer multi-qubit quantum circuits. Shin et al. (2023) proposed an encoding scheme that enables exponential growth of the spectrum of the data re-uploading circuit. On this basis we consider the coefficients of the data re-uploading circuit and analyze them, thus demonstrating that under optimal conditions data re-uploading circuits have an exponential advantage over classical methods in fitting the Fourier series. Meanwhile, our approach is more practical, unlike the above studies which only perform experiments to verify mathematical properties, we conducted experiments on real-world datasets and establish a bridge between data re-uploading circuits and implicit neural representations.
@@ 4.3 Theoretical Analysis of QIREN
By observing Eq. (13), we can understand that applying a Linear layer before uploading the data to the quantum circuit is equivalent to adjusting the eigenvalues in the Hamiltonian, ultimately affecting the frequency spectrum. Using the Linear layer brings two advantages. First, it can make the frequency spectrum larger. From Eq. (18) and (19), it can be deduced that certain sums of eigenvalues generate the same term in the spectrum. This redundancy can be reduced by using the Linear layer, and the size of the spectrum can be extended from $(2dL+1)^{d_{h}}$ to up to $((3^{d}-1)L+1)^{d_{h}}$. Second, it enables the frequency spectrum to be adjustable, aiming to cover frequencies with larger coefficients. Therefore, incorporating Linear layers can further enhance the fitting performance of QIREN. A more detailed explanation of Claim 2 can be found in Appendix B.
@@ Appendix B Additional Explanation on Claim 2
The role of the Linear layer in QIREN is actually an improved version of the exponential encoding scheme (Shin et al., 2023). We will next explain the role of the Linear layer and conclude with a comparison with the exponential encoding scheme that illustrates the advantages of the Linear layer encoding scheme.
If we ignore the bias term, the Linear layer essentially maps the input $x$ to $(w_{1}x,\ldots,w_{d}x)$. Considering the form of the encoding gate as $S(x)=\otimes_{q=1}^{d}\mathrm{e}^{-\mathrm{i}(w_{q}x)Z/2}=\otimes_{q=1}^{d}\mathrm{e}^{-\mathrm{i}x(w_{q}Z/2)}$, the specific form of $H$ is as follows:
$H=\sum_{q=1}^{d}w_{q}Z^{(q)}/2.$ (A3)
In the exponential encoding scheme, $H=\sum_{q=1}^{d}\beta_{q}Z^{(q)}/2.$ The difference between the two methods is that $w_{q}$ is a learnable parameter while $\beta_{q}$ is a hyperparameter that needs to be manually tuned. In the absence of degeneracy, $H$ can have $2^{d}$ distinct eigenvalues. Thus, the spectrum can contain at most $14^{d}-2^{d}+$ 1 distinct frequencies (including zero). But if in the absence of the Linear layers i.e. $w_{q}=1$, then there will be only $d+1$ distinct eigenvalues and the size of the spectrum $\Omega=\left\{\lambda_{k}-\lambda_{j}\right\}$ will be limited to $12d+$ 1. Next we derive a theoretical upper bound on the spectrum that can be extended by the Linear layer, where the derivation follows the idea in the exponential encoding scheme (Shin et al., 2023).
@@
With $H=w_{1}Z/2$, we have $\Omega^{(1)}=\{-w_{1},0,w_{1}\}$ as the frequency spectrum. If we append one more encoding gate to $H$,we have $\Omega^{(2)}=\{-w_{2}-w_{1},-w_{2},-w_{2}+w_{1},-w_{1},0,w_{1},w_{2}-w_{1},w_{2},w_{2}+w_{1}\}$. More generally, let $\Omega^{(k)}$ as the $k$-th frequency spectrum as a result of using $k$ encoding gates. Then $\Omega^{(k)}$ has all elements from $\Omega^{(k-1)}$, along with the new ones generated by adding ±$\pm w_{k}$ to all elements of $\Omega^{(k-1)}$:
$\Omega^{(k)}=\left\{\Omega^{(k-1)}-w_{k},\Omega^{(k-1)},\Omega^{(k-1)}+w_{k}\right\}.$ (A4)
@@
As $\Omega^{(k)}$ is always symmetric about zero, to generate a maximally non-degenerate frequency spectrum, the inequality
$\max\left\{\alpha\in\Omega^{(k-1)}\right\}<w_{k}-\max\left\{\alpha\in\Omega^{(k-1)}\right\}$ (A5)
@@
is to be satisfied, where we can easily deduce that $\max\left\{\alpha\in\Omega^{(k-1)}\right\}=\sum_{j=1}^{k-1}w_{j}$. When Eq. (A5) is satisfied, the spectrum can be extended from $12d+$ 1 to the theoretical upper bound $3^{d}$.
Finally, we compare the exponential encoding scheme (Shin et al., 2023) with our Linear layer encoding scheme. First, exponential encoding scheme can always reach the theoretical upper bound of the extension of the spectrum by manually selecting the hyperparameters. However, when faced with signals with different spectrum, manual tuning obviously lacks flexibility. Second, after the neural network layers are stacked, the spectrum of the function to be fitted in each layer is unknown and the optimal hyperparameters cannot be manually selected. Our Linear layer encoding scheme, on the other hand, is learnable and utilizes gradient descent to achieve a suitable trade-off between reducing spectrum redundancy and covering the signal frequency. This allows us to implement an end-to-end training process without the need to manually tune the hyperparameters. Thus our approach is more practical and can be applied to complex real-world tasks.